OpenStudy (anonymous):

The ratio of boys to girls in a classroom is 2 to 7. If there are 63 students in the room, how many are girls ?

OpenStudy (anonymous):

let the number of boys be 2x let the number of girls be 7x 2x+7x=63 9x=63 x=63/9 x=7 girls=7x=7*7=49 No of girls are 49 hope this helped :)

OpenStudy (jack1):

props @candiee , perfect explanation hey

OpenStudy (anonymous):

tnksss @Jack1

OpenStudy (anonymous):

Thank you so much @ candiee

OpenStudy (anonymous):

@candiee